### v1

# Understanding differences in applying DETR to natural and medical images

Abstract
Transformer-based detectors have shown success in computer vision tasks with natural images. These models, exemplified by the Deformable DETR, are optimized through complex engineering strategies tailored to the typical characteristics of natural scenes. However, medical imaging data presents unique challenges such as extremely large image sizes, fewer and smaller regions of interest, and object classes which can be differentiated only through subtle differences. This study evaluates the applicability of these transformer-based design choices when applied to a screening mammography dataset that represents these distinct medical imaging data characteristics. Our analysis reveals that common design choices from the natural image domain, such as complex encoder architectures, multi-scale feature fusion, query initialization, and iterative bounding box refinement, do not improve and sometimes even impair object detection performance in medical imaging. In contrast, simpler and shallower architectures often achieve equal or superior results. This finding suggests that the adaptation of transformer models for medical imaging data requires a reevaluation of standard practices, potentially leading to more efficient and specialized frameworks for medical diagnosis.
Keywords: Machine Learning, Vision Transformers, Object Detection, Breast Cancer
1 Introduction
Figure 1: An overview of the study. In this work, we examine design choices in Deformable DETR. To evaluate them, we used the NYU Breast Cancer Screening Dataset. Specifically, we assess six design choices (highlighted in red): 1. input resolution; 2. number of encoder layers; 3. whether to use multi-layer feature fusion; 4. number of object queries; 5. query initialization method; 6. whether to use the iterative bounding box refinement method. The graph on the right demonstrates the performance changes, indicated by Average Precision (AP), after applying these design choices. +(2): reducing the number of encoder layers; +(3): removing multi-layer feature fusion; +(4): reducing the number of object queries; +(5): adding object query initialization method; +(6): adding iterative bounding box refinement.) The results reveal that a simplified architecture +(2,3,4) is more suitable for medical datasets and results in improved performance. Increasing complexity of the model, as indicated by +(2,3,4,5,6), brings marginal benefit.
Recent advancements in computer vision have increasingly focused on leveraging transformer architectures (Vaswani et al., 2017) for image classification and object detection (Dosovitskiy et al., 2020; Liu et al., 2021; Carion et al., 2020; Touvron et al., 2021). Thanks to their inherent self-attention mechanisms, transformers effectively capture global dependencies and understand contextual relations across entire images, enabling them to acquire more expressive features. Owing to these advantages, transformer-based models have become prevalent in natural imaging analysis. Moreover, their application in medical imaging has shown promising results, indicating their potential to transform this field as well. (Chen et al., 2021; Dai et al., 2021b; Valanarasu et al., 2021; Zheng et al., 2022).
Object detection plays a crucial role in medical image analysis as detection models directly identify the locations of abnormalities, which is critically important in application of AI in medical diagnosis. With the rise of transformer-based models, Detection Transformer (DETR), an end-to-end transformer-based object detection framework (Carion et al., 2020) has become popular. DETR does not require Non-Maximum Suppression (NMS) (Girshick et al., 2014), a post-processing step that reduces overlapping detections, which is non-differentiable and sensitive to predefined thresholds. The elimination of NMS enables end-to-end training and direct optimization of the objective function. DETR has established itself as one of the most competitive detection framework by consistently achieving state-of-the-art accuracy on MS-COCO datasets (Zhu et al., 2020; Zhang et al., 2022; Zong et al., 2023). Consequently, the potential of DETR draw significant research interest in improving overall performance and accelerating learning (Zhu et al., 2020; Zhang et al., 2022; Chen et al., 2022b; Wang et al., 2022; Chen et al., 2022a). This has led to the development of some highly engineered DETR-based architectures and a diverse range of design choices within the DETR family.
Despite the success of DETR architectures for natural image data, their application to medical imaging poses challenges. Medical images differ significantly from natural images in several aspects:
  1. 1.
    High resolution and small regions of interest Medical images are usually in very high resolutions with small, critical regions of interest (e.g., lesions, masses and calcifications) (Moawad et al., 2023; Heath et al., 2001).
  2. 2.
    Standardized acquisition Medical images are acquired through very standardized procedures, resulting in images consisting of three elements: the body part of interest, the surrounding body tissue, and black pixels outside the body. In contrast, the backgrounds in natural images can contain anything existing in the natural world.
  3. 3.
    Small number of objects per image Medical images typically focus on a narrow range of abnormalities, resulting in fewer objects of interest within each image from very few classes.
  4. 4.
    Small and imbalanced data sets Medical imaging data sets are often smaller and exhibit a more imbalanced class distribution, as positive cases (i.e. unhealthy subjects) are usually a lot less common than negative cases (i.e. healthy subjects)(Galdran et al., 2021; Heath et al., 2001; Wang et al., 2017).
Many highly engineered design choices of DETR-family models, such as multi-scale feature fusion, iterative bounding box refinement, works well for natural image data. However, when applying to medical images, where data characteristics differ markedly, we need to reexamine the effectiveness of those engineering design choices. Based on the differences between medical and natural images, we hypothesize that those design choices make DETR architecture unnecessarily complex, increasing its computation demand and memory consumption, and might not even work well with the unique properties of medical imaging data listed above. We hypothesize that a simplified model, tailored to the characteristics of medical imaging data, might be more efficient and accurate.
To verify the aforementioned hypothesis, this work investigates whether design choices in the DETR family, originally intended for natural images, remain beneficial or even necessary for medical imaging applications. We use the Deformable DETR as our baseline model and the NYU Breast Cancer Screening Dataset as a representative case study (Wu et al., 2019). This dataset contains high-resolution screening mammography with small lesions, serving as a representative dataset for the characteristics of medical imaging mentioned above. The key findings of our work are:
  • Models with a reduced number of encoder layers and no multi-scale feature fusion learn faster without compromising detection performance. These changes actually increased performance by $1.1\%$ in $\mathrm{AP}_{10,50}$, while accelerating the training by 50%.
  • Additional techniques applied in the decoder, such as object query initialization methods and iterative bounding box refinement, do not improve detection performance on medical image datasets as they do on natural image datasets. Instead, they resulted in a 0.7% drop in $\mathrm{AP}_{10,50}$, indicating these techniques may not be suitable for medical imaging tasks.
  • Increasing the number of object queries improves detection performance up to a point. However, unlike for natural image data, using an excessive number of queries (e.g., 800) leads to a 2% drop in performance due to an increase in false positives.
  • Design choices aimed at localizing ROI candidates, such as object query initialization methods, iterative bounding box refinement and more object queries, do not improve performance, suggesting that correctly classifying ROI candidates is key to performance improvement.
2 Background on DETRs
DETR offers several advantages over commonly applied detection models such as Mask-RCNN (He et al., 2017) and YOLO (Redmon et al., 2016). Its transformer-based architecture enables learning more expressive features than CNN-based models, and end-to-end training improves optimization and performance. However, DETR suffers from slow learning. To address this issue, several studies have proposed techniques to accelerate learning and enhance overall performance (Zhu et al., 2020; Wang et al., 2022; Chen et al., 2022b; Zhang et al., 2022). Deformable DETR stands out for its competitive performance on the MS COCO dataset (Lin et al., 2014). It introduces a deformable attention module, which reduces training time by a factor of 10 and enables multi-scale feature fusion that improve detection, especially for small objects. Given its widespread adoption as a benchmark in the subsequent research works (Roh et al., 2021; Dai et al., 2021a; Zhang et al., 2022; Yao et al., 2021), we choose Deformable DETR as the baseline architecture for our experiments. This section details the essential components of DETR and Deformable DETR architectures.
DETR
DETR consists of a backbone, an encoder-decoder transformer, and a prediction head, as illustrated in Figure 2(a).
Given an input image $x\in\mathbb{R}^{C_{0}\times W_{0}\times H_{0}}$, the backbone network $f$ produces a low-resolution activation map $x_{s}=f(x)\in\mathbb{R}^{C\times W\times H}$. This map is further processed by a $1\times 1$ convolution to collapse the channel dimension $C$ into a smaller size $d$, resulting in image tokens $x_{f}=\mathrm{conv}(x_{s})\in\mathbb{R}^{WH\times d}$. To preserve spatial information in the original image, each token is paired with a positional encoding, denoted by $x_{p}\in\mathbb{R}^{WH\times d}$. The encoder is a standard attention-based transformer where each layer consists of a multi-head self-attention module (MHSA) followed by a feedforward network (FFN). For an in-depth formalization of MHSA, refer to Appendix A.1. Typically, the DETR encoder consists of 6 layers. The encoder preserves the dimension of the input, producing $x_{enc}\in\mathbb{R}^{WH\times d}$.
The decoder receives two inputs, the encoded features $x_{enc}$ and $N$ object queries $q\in\mathbb{R}^{N\times d}$. Object queries play a central role in DETR architecture. They are learnable embeddings that work as placeholders for the potential objects in an image. Each of them attends to the specific regions of the image and is individually decoded into a bounding box prediction. Each object query is the sum of two learnable embedding: content embeddings $q_{c}\in\mathbb{R}^{N\times d}$, initialized as zero vectors, and the positional embeddings $q_{p}\in\mathbb{R}^{N\times d}$, indicating each query’s position. More methods for initializing object queries are discussed in Section 3. Decoder layers consists of a MHSA, enabling inter-query learning, and multi-head (MH) cross-attention to integrate encoder features, and a FFN. The formalization of MH cross-attention is detailed in Appendix A.2.
After the decoder, each object query is independently decoded into bounding box coordinates and class scores through a three-layer FFN and a linear layer respectively.
(a) vanilla DETR
(b) Deformable DETR
Figure 2: Architecture of DETR and Deformable DETR. Both architectures consist of a backbone, an encoder-decoder transformer and a prediction head. Deformable DETR differs from DETR in its utilization of multiple feature level fusions and its application of deformable attention instead of standard attention mechanism.
Deformable DETR
Deformable DETR improves upon DETR by introducing a deformable attention module, which accelerates training and enhances the detection of small objects. The architecture of Deformable DETR is illustrated in Figure 2(b).
Unlike the standard attention mechanism that calculates attention scores between all query-key pairs, resulting in $(WH)^{2}$ pairs for a feature map of size $W\times H$, deformable attention selectively computes attention scores on a subset of $k<<WH$ keys for each query. The subset is selected through a learnable key sampling function, allowing the model to focus on the most informative regions for each query. For a detailed formalization of the deformable attention module, refer to Appendix B.1.
For dense prediction tasks such as object detection, incorporating higher-resolution feature maps can substantially improve detection performance, especially for smaller objects (He et al., 2017). However, the complexity of the standard attention mechanism is quadratic with respect to the number of tokens, making it infeasible for multiple scales of feature maps. The deformable attention mechanism enables effective multi-scale feature fusion. Specifically, the encoder receives the output feature maps $x_{1},x_{2},x_{3}$ from the backbone’s last three layers, and a convolutional layer generates the lowest resolution feature map $x_{4}$. All four feature maps undergo a $1\times 1$ convolution and then are reshaped into a sequence of feature vectors of dimension $d$, denoted by $x_{f}\in\mathbb{R}^{M\times d}$. Each token is associated with a positional embedding, as well as a layer embedding to identify feature map level. Section 3 explores the benefits of multi-feature fusion for medical imaging datasets.
Moreover, Deformable DETR introduces reference points in the deformable attention module. In the encoder, each query $q$ is associated with a 2D reference point $p_{q}=[x,y]$, denoting its location on the feature map. The key sampling function generates $k$ sampling offsets with respect to the reference point, and thus determines the $k$ keys for the query. Similarly in the decoder, the reference point of each object query $q$ is defined by a linear projection of its positional embedding $q_{p}$. In this way, each object query can be mapped to a position on the feature map. This approach allows object queries to focus on specific regions, significantly accelerating learning Zhu et al. (2020).
DETR in Medical Imaging
DETR-based architectures have been widely applied to various medical imaging tasks, often with architectural tweaks to improve overall performance. For example, Mathai et al. (2022) leveraged a bounding box fusion technique on DETR to reduce false-positive rate in lymph nodes detection. MyopiaDETR (Li et al., 2023) utilizes a Feature Pyramid Network to improve detection of small objects in lesion detection of pathological myopia. COTR (Shen et al., 2021) embeds convolutional layers into DETR encoders to accelerate learning in polyp detection. Although these works achieved good performances, our experiments indicates that, contrary to the common understanding, simplifying the DETR architecture can improve accuracy and accelerate training. One work we identified that also points in this direction, Cell-DETR (Prangemeier et al., 2020), also reduces the number of parameters by ten times, achieving faster inference speeds while maintaining performance on par with state-of-the-art baselines. Finally, Garrucho et al. (2023) applied out-of-the-box Deformable DETR on mammography for mass detection. However, their focus is the effect of a data augmentation method on its detection performance. Despite these advances, a systematic exploration of the effectiveness and relevance of foundational DETR design choices remains under-explored.
3 Methods
3.1 Design Choices
In this section, we outline key design choices of Deformable DETR relevant to the unique characteristics of medical images: input resolution, the number of encoder layers, multi-scale feature fusion, the number of object queries, and two techniques enhancing the decoding process, query initialization and iterative bounding box refinement (IBBR). We investigate if incorporating these elements into model design yields improvements in detection performance on medical imaging data to the same degree they do on natural image datasets.
Input resolutions
Downsampling input images is a standard practice in detection models for computational efficiency and to satisfy memory constraints. Natural images can be significantly downsized without losing important features such as edges, shapes, and textures necessary for accurate predictions. Data from many medical imaging modalities, such as X-ray images, CT scans, and whole slide images, are at least an order of magnitude larger. These high-resolution medical images contain fine-grained details such as small lesions or slight changes in tissue density that are crucial for accurate diagnosis (Sabottke and Spieler, 2020; Thambawita et al., 2021). On the other hand, processing high-resolution medical images without any downsampling is often infeasible due to the high computational requirements. Hence, we conduct experiments on medical images with resolution ratios ranging from 25% to 100% to examine the optimal input resolution that balances model accuracy with computational efficiency and memory usage.
Encoder complexity
Medical image datasets have three distinct characteristics. Firstly, they are typically smaller compared to natural image datasets due to the limited number of patients with specific conditions. Secondly, medical images within one dataset are usually very homogeneous, depicting single body part such as the brain, breast, or chest, with uniform textures and grayscale. Furthermore, while natural images contain hundreds or thousands of objects classes, medical image datasets usually have much fewer classes. For example, NIH Chest X-ray (14 disease classes) (Wang et al., 2017), DDSM (Heath et al., 2001) (2 classes), BraTs (4 classes) (Moawad et al., 2023). This means that, overall, there is much less variation in the data that the network has to model. Given the broadly accepted idea that model complexity should match task complexity (Geman et al., 1992), we suspect that simpler, shallower architectures might be better suited to these medical imaging datasets to mitigate overfitting and enhance training efficiency. Additionally, objects in medical images tend to have more uniform sizes compared to those in natural images. For instance, the standard deviation of standardized object sizes [^1] Standardized object size refers to the area of the bounding box of an object divided by the area of the image, providing a normalized measure of object size relative to the image size. in the MS COCO dataset is 0.16, whereas it is only 0.025 in the NYU Breast Cancer Screening Dataset. This observation leads us to question the necessity of employing multi-scale feature fusion for medical images, which is more beneficial for detecting objects of varying sizes. To explore these hypotheses, we experimented with modifications to the encoder of Deformable DETR, including reducing the number of encoder layers and utilizing fewer scales of feature maps from the backbone.
Number of object queries
In DETR, each object query is individually decoded into a bounding box prediction. Thus, the total number of object queries determines the number of bounding boxes predicted per image (Carion et al., 2020). Most DETR models are optimized for natural image datasets such as MS COCO, where a single image can contain up to $100$ objects. Consequently, the number of object queries is usually set to $300$ in DETR models. However, a single medical image typically contains fewer than $10$ objects, with the majority containing only $1$ or no objects at all. For this reason, the hyperparameters suggested in DETR research might not be suitable for medical images, making it crucial to adjust this number for medical image datasets.
Decoding techniques
Object queries initialization and iterative bounding box refinement (IBBR) are two techniques employed by many DETR models to facilitate the decoding process of the object queries (Zhu et al., 2020; Zhang et al., 2022; Yao et al., 2021). These methods have proven effective in boosting detection performance on natural image datasets, increasing AP by 2.4 on the MS COCO dataset (Zhu et al., 2020). This study aims to evaluate their suitability for medical data. We tested three initialization strategies for the positional and content embeddings of object queries, as characterized by Zhang et al. (2022).
  • Static queries: Both positional and content embeddings are initialized as random learnable embeddings. This generic starting point requires the model to learn from scratch where objects are likely located and what features represent those objects. While offering maximum flexibility, this method may lead to slower learning as the model has to discover spatial and content relationships without prior knowledge. The standard Deformable DETR uses this approach.
  • Pure query selection: Both content and positional embeddings are initialized with selected encoder features. To choose $K$ encoder features that most likely contain objects, the model either uses a regional proposal network (Yao et al., 2021; Chen et al., 2022b) or applies the prediction head to encoder features (Zhu et al., 2020). Our experiments employ the latter approach. This method leverages existing knowledge encoded in the encoder features, guiding the model more directly toward accurate object detection and significantly speeding up the learning process.
  • Mixed query selection: This method combines the previous two methods. By using pure query selection for positional embeddings and static embeddings for content, the model is informed about the likely positions of objects through the spatial priors while retaining flexibility in learning content representations from scratch. DINO (Zhang et al., 2022) suggests that mixed query selection methods provide the best model performance among all three methods.
IBBR, first introduced in Deformable DETR, iteratively updates the reference points of object queries towards the objects of interest in each image. Reference points guide the model’s deformable attention mechanism, indicating which regions to search for objects. These reference points are distributed randomly across the image, ensuring a broad coverage without any prior knowledge about where objects might be located. With IBBR, these reference points can move progressively towards the objects through each decoder layer, providing more accurate signals for the attention. This technique has been applied extensively in DETR family (Zhu et al., 2020; Chen et al., 2022b; Wang et al., 2022; Liu et al., 2022) and has shown to effectively speed up training and improve detection performance. Detailed formalization of this method is in Appendix C.
3.2 Data and task
We conduct experiments on the NYU Breast Cancer Screening Dataset (Wu et al., 2019), representative for the characteristics of medical image datasets. This dataset contains $229,426$ digital screening mammography exams from $141,472$ patients screened at NYU Langone Health. Each exam includes a minimum of four images, each with a resolution of $2944\times 1920$, covering two standard screening views: craniocaudal (CC) and mediolateral oblique (MLO), for both the left and right breasts. An example of a mammography exam is shown in Figure 3. The dataset is annotated with breast-level cancer labels indicating biopsy-confirmed benign or malignant findings. Moreover, the dataset also provides bounding box annotations, and class labels (benign or malignant) of each visible positive findings. The entire dataset contains $985$ breasts with malignant findings and $5,556$ breasts with benign findings. The dataset is divided into training ($82\%$), validation ($5\%$) and test sets ($13\%$) ensuring a proportional distribution of both benign and malignant cases across the subsets.
Figure 3: An example screening mammography exam. From left to right: left craniocaudal view, left mediolateral oblique view, right craniocaudal view, right mediolateral oblique view. This patient has a benign lesion in the left breast. It is marked with a red bounding box from both views of the left breast.
3.3 Evaluation Metrics
In this study, we focus on evaluating the models’ ability to detect malignant lesions. We use Average Precision (AP) (Everingham et al., 2010) and the Free-Response Receiver Operating Characteristic curve area (FAUC) (Bandos et al., 2009), which is a frequently used metric in medical image analysis (Yu et al., 2022; Wang et al., 2018; Petrick et al., 2013). Specifically, we focus on FAUC at the rate of 1 false positive per image, referred to as $\text{FAUC}^{1}$, in line with the approach described by Bandos et al. (2009). Following prior works (Jailin et al., 2023; Kolchev et al., 2022; Konz et al., 2023), we define a positive bounding box to have at least 10% 50% Intersection over Union (IoU) with a ground truth box. These thresholds are deemed more appropriate for accurately detecting small-sized objects, such as cancerous lesions. $\mathrm{FAUC}^{1}$ at 0.1 Iou threhold is denoted as $\text{FAUC}^{1}_{10}$. Following the notation of COCO evaluation metrics (Lin et al., 2014), we denote AP at 0.1 IoU threshold as $\text{AP}_{10}$. Additionally, we report the average AP across IoU thresholds ranging from 0.1 to 0.5, in steps size of 0.05, denoted as $\text{AP}_{10,50}$.
To clearly explain how well our models detect objects, we differentiate between “localization” and “classification.”
  • Localization refers to the task of accurately drawing a bounding box around each ground-truth object. To be consistent with the definition of FAUC and AP, an object is considered successfully localized if the model produces a bounding box overlapping the ground truth box by more than 10% IoU. To quantify localization accuracy, we compute the percentage of ground-truth objects successfully detected by the model. Assume there are $m$ ground-truth objects $G_{i}$ where $i=1,\ldots,m$ and $p$ predicted boxes $P_{j}$ where $j=1,\ldots,p$ in an image. The maximum IoU for a ground-truth bound box $G_{i}$ among all the predicted bounding boxes $P_{j}$ is $\max_{j}(\mathrm{IoU}(P_{j},G_{i}))$. Localization accuracy $L$ is then expressed as
    $L=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j}\left({\mathrm{IoU}\left(P_{j},G_{i}\right)}\right)\geq 0.1}.$ (1)
  • Classification involves associating the object inside each predicted box with the correct class. We consider models’ classification accuracy using the percentage of successfully localized objects among the predicted bounding boxes with the top 10 highest predicted scores in each image. Among all the predicted bounding boxes $P_{j}$ in an image, let $S$ be the subset of indices of the top 10 predicted bounding boxes in an image. Localization performance considering classification is expressed as
    $L_{\mathrm{top10}}=\frac{1}{m}\sum^{m}_{i=1}\mathds{1}_{\max_{j\in S}({\mathrm{IoU}(P_{j},G_{i})})\geq 0.1}.$ (2)
3.4 Experimental Setup
Our baseline model is a Deformable DETR in its default setting, using a Swin-T Transformer (Liu et al., 2021) as the backbone. The backbone is pretrained with a breast cancer classification task on the same dataset (see Appendix D for details). All models are optimized using the AdamW optimizer (Loshchilov and Hutter, 2017) with a step scheduler for $60$ epochs. The step scheduler reduces the learning rate by a factor of $0.1$ for the last $20$ epochs. We tuned the remaining hyperparameters using random search as detailed in Appendix E. For each experiment, we trained $5$ models with different random seeds and we reported their performance mean and standard deviation.
4 Results
Our experiments with the five design choices, including input resolutions, encoder layer complexity, multi-scale feature fusion, number of object queries, and two decoding techniques, all demonstrate that modifications of the standard Deformable DETR do not align well with the unique characteristics of medical imaging datasets. This misalignment results in unnecessary computational overhead and sub-optimal performance.
Input Resolution
Our experiments reveal a positive correlation between input resolution and detection performance up to a certain threshold (Table 1). Specifically, increasing the resolution ratio from a quarter size ($0.25$) to half size ($0.5$) significantly improves the performance, with a $9.8\%$ increase in $\text{FAUC}^{1}_{10}$, $8.6\%$ in $\text{AP}_{10}$ and $6.4\%$ in $\text{AP}_{10,50}$. Further increasing the resolution to $0.75$ continues to boost performance, though with a smaller margin. However, using full-resolution images results in a performance drop across all metrics. In high-resolution images, objects of interest can have defining features spread across a large area, requiring the model to integrate clues from a broader area. The deformable attention mechanism only focuses on a selective set of keys centered around the reference points, which may miss necessary information in high resolution images. This phenomenon of diminished performance at very high resolutions has been observed by prior studies (Sabottke and Spieler, 2020; Thambawita et al., 2021). Richter et al. (2021) also challenged the notion that higher resolution always equates to better performance and pointed out that different CNN networks have a preferred input size.
It is also noteworthy that the increase in resolution from quarter to full resolution comes at the cost of ×$\times 15$ increase in GFLOPs. To balance the training efficiency and detection performance, subsequent experiments are conducted for half-resolution images.
Table 1: Standard Deformable DETR performance using different input image resolutions. The full image resolution is $2944\times 1920$. The detection performance is measured by AP and FAUC, as defined in Section 3.3. GFLOPs are reported using one billion floating point operations per second as the unit. Using a 75% scale of full resolution yields the best detection performance.
Image Resolution $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$ GFLOPs
$1.0$ $0.676\pm 0.010$ $0.671\pm 0.005$ $0.477\pm 0.006$ $4367$
$0.75$ $0.688\pm 0.010$ $0.683\pm 0.007$ $0.490\pm 0.007$ $2448$
$0.5$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $1102$
$0.25$ $0.591\pm 0.016$ $0.583\pm 0.007$ $0.400\pm 0.021$ $284$
Encoder complexity: number of encoder layers
In this set of experiments, we examined the impact of the number of encoder layers and the necessity of muilti-scale feature fusion in the Deformable DETR architecture.
Table 2: Varying the number of encoder layers in Deformable DETR. The standard Deformable DETR has $6$ encoder layers. For both backbones, we do not observe any significant performance drop when using fewer encoder layers. A small performance decrease is observed for a model with no encoder at all.
backbone #encoder $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$params GFLOPs
layers
ResNet50 00 $0.643\pm 0.020$ $0.619\pm 0.010$ $0.421\pm 0.025$ $35.4$ $536$
ResNet50 $1$ $0.656\pm 0.011$ $0.624\pm 0.016$ $0.439\pm 0.005$ $36.2$ $624$
ResNet50 $3$ $0.655\pm 0.016$ $0.627\pm 0.013$ $0.439\pm 0.015$ $37.7$ $801$
ResNet50 $6$ $0.657\pm 0.009$ $0.626\pm 0.009$ $0.436\pm 0.011$ $40.0$ $1067$
Swin-T 00 $0.681\pm 0.013$ $0.662\pm 0.013$ $0.458\pm 0.023$ $35.9$ $570$
Swin-T $1$ $0.684\pm 0.009$ $0.672\pm 0.011$ $0.463\pm 0.014$ $36.7$ $659$
Swin-T $3$ $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $38.3$ $836$
Swin-T $6$ $0.689\pm 0.008$ $0.669\pm 0.012$ $0.464\pm 0.019$ $40.5$ $1102$
As the backbone critically affects feature extraction, to reach more generalizable conclusions, we conducted this experiment with two distinct backbones, ResNet50 and Swin-T. For both backbones, models with merely $1$ and $3$ encoder layers matched the performance of models with $6$ encoder layers across all three detection metrics (Table 2). Impressively, this reduction in the number of layers resulted in up to $40\%$ decrease in GFLOPs. Furthermore, when using Swin-T as the backbone, the encoder-free models achieve nearly the same detection performance as their 6-layer counterparts, with less than a $1\%$ drop across all three detection metrics, while cutting the GFLOPs almost by half.
This observation aligns with the recent development of the encoder-free $\text{D}^{2}\text{ETR}$ (Lin et al., 2022), which outperforms the standard DETR model on the MS COCO dataset (Lin et al., 2014). Such findings challenge the traditional view that encoders are essential for feature transformation and multi-level feature integration within DETR models. Our results suggest that it is possible to create effective DETR architectures without encoders, avoiding significant sacrifices in performance. Additionally, a strong backbone can further reduce DETR’s reliance on the encoders, further supporting the potential for more streamlined designs.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 4: (a) Comparison of the number of objects per image and object sizes between MS COCO and the NYU Breast Screening dataset. Compared to the COCO dataset, in the NYU Dataset object sizes vary less. The standard deviation of object sizes in COCO and NYU breast are 0.161 and 0.025 respectively. (b) An example image from the NYU Breast Dataset containing a cancerous lesion. In this data set, there are, on average, 1.10 objects per image in images that contain any objects. (c) Example images from the MS COCO dataset. Both images contain multiple objects in different sizes and some of them even overlap, adding difficulty to object detection. There are, on average, 7.33 objects per image in MS COCO dataset.
Encoder complexity: Multi-scale feature fusion
Standard Deformable DETR applies the attention mechanism to four scales of feature maps in the encoder: three from the last three layers of the backbone and fourth from a convolution layer applied to final output of the backbone, as illustrated in Figure 2(b). Previous findings highlight that using multi-scale feature fusion improves detection performance by 1.3 on the MS COCO dataset and also on other datasets (He et al., 2017; Zhou et al., 2021; Zeng et al., 2022), our results in Table 3 indicate that comparable performance can be achieved using only the output of the last layer of backbone. This suggests that multi-scale feature fusion may not be crucial for detecting abnormalities in medical images.
The characteristics of our dataset likely explain this finding. Firstly, our dataset predominantly consists of images containing a single object. Secondly, the sizes of these objects are relatively uniform. These characteristics contrast with natural image datasets such as MS COCO, where object sizes vary widely due to perspective, distance from the camera, and the inherent size differences between object classes, as shown in Figure 4(c). Multi-scale feature fusion allows models to be sensitive to this size variability by utilizing feature maps at different scales, each of which captures details at different resolutions. However, our dataset is more homogeneous, with images containing single objects of similar sizes (Figure 4(b). For such data, the benefits of multi-scale feature fusion are less pronounced. Figure 4(a) contrasts the object size distribution between the two datasets. In our dataset, the majority of images contain only a single object and most objects take only approximately $1\%$ of the entire image, whereas MS COCO dataset exhibits a broader variation in both the sizes of objects and the number of objects per image. Consequently, in a homogeneous dataset, the additional complexity of multi-scale feature fusion may not translate into better performance.
Table 3: The performance of Deformable DETRs using different combinations of feature levels. The standard Deformable DETR uses the 4 levels of feature maps from the backbone (three from the backbone and the fourth layer from a convolution layer applied to final output of the backbone). Using only the 3rd feature level, the last layer of the backbone, achieves on-par performance with using multi-level feature fusion. All experiments are conducted on model with 3 encoder layers (cf. Table 2).
Feature Levels $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$ params GFLOPs
$1,2,3,4$ (standard) $0.688\pm 0.012$ $0.677\pm 0.011$ $0.470\pm 0.019$ $38.3$ $836$
$1$ $0.637\pm 0.012$ $0.627\pm 0.009$ $0.434\pm 0.009$ $35.5$ $734$
$2$ $0.680\pm 0.005$ $0.666\pm 0.009$ $0.467\pm 0.016$ $35.6$ $570$
$3$ $0.688\pm 0.004$ $0.675\pm 0.007$ $0.475\pm 0.006$ $35.7$ $528$
Number of object queries
(a) $\text{FAUC}^{1}_{10}$
(b) $\text{AP}_{10}$
(c) $\text{AP}_{10,50}$
(d) $L$ and $L_{top10}$
Figure 5: (a)-(c) The detection performance of Deformable DETR with varying number of object queries. Detection performance improves as the number of object queries increases from 5 to 100. However, performance declines when the number exceeds 100. (d) The localization performance $L$ continues to improve as the number of object query increases. The classification performance $L_{\mathrm{top10}}$ worsens as number of object queries surpass 100, indicating an increase in false positives.
Figure 5(a)-(c) depicts the impact of increasing the number of object queries from 5 to 500 on three detection metrics. Increasing the number of object queries from $5$ to $100$ consistently improves detection performance. However, exceeding this number does not bring additional benefits and may even slightly hurt performance. The benefit of more object queries for object localization is evident in Figure 5(d), which shows improved localization performance $L$ (as defined in Equation 1) as the number of queries increases. However, better localization does not necessarily equate to improved overall detection performance, as indicated by $L_{\mathrm{top10}}$ (cf. Equation 2), the proportion of correctly localized objects that are also among top $10$ classification scores. Although the model correctly identifies more ground truth objects, it assigns them low classification scores, resulting in fewer objects ranked among the top 10 bounding boxes.
We hypothesize that having more object queries increases the chances of localizing false positives. More object queries expand model’s search space, making it more sensitive to variations in subtle patterns or textures that may resemble the characteristics of positive objects. These false positives cause true positives to be ranked lower. In summary, while more object queries improves the chances of detecting an object, they also introduce the risk of misidentifying irrelevant entities as objects of interest, thus adversely impacting the model’s overall performance. This phenomenon can be more severe in medical imaging datasets, where there are significantly fewer objects (usually a single object) within an image, increasing the likelihood of false positives overwhelming the true positives.
Decoding Techniques
We implemented two techniques commonly adopted in DETR family using the simplified model with $3$ encoder layers and $1$ feature map: query initialization method and iterative bounding box refinement (IBBF). Surprisingly, empirical results indicate that neither technique significantly improved detection performance across all three metrics, as detailed in Table 4. Following previous analysis, we examine the model’s localization and classification performance separately by evaluating on $L$ and $L_{\mathrm{top10}}$. We found that while these techniques enhanced localization performance, they adversely affected classification performance, as depicted in Figure 6.
We visualize training and validation losses for localization and classification separately in Figure 7. Localization losses include the IoU loss and bounding boxes regression loss, while classification loss is the BCE loss on all prediction boxes. The figure reveals that the model equipped with either decoding techniques tends to overfit the training data more than the model without them, particularly in classification loss, as illustrated in Figure 7. This overfitting is likely the reason behind the observed decline in classification performance. We hypothesize that this is due to the significantly smaller number of positive objects in our dataset.
Table 4: Deformable DETR with different query initialization methods and IBBF. Neither of the techniques improve detection performance across all three metrics. All experiments are conducted on model with $3$ encoder layers and without multi-scale feature fusion technique.
Refinement Query Initial. $\text{FAUC}^{1}_{10}$ $\text{AP}_{10}$ $\text{AP}_{10,50}$
Static $0.688\pm 0.004$ $0.677\pm 0.011$ $0.470\pm 0.019$
Pure $0.678\pm 0.007$ $0.668\pm 0.011$ $0.474\pm 0.018$
Mixed $0.670\pm 0.012$ $0.633\pm 0.010$ $0.443\pm 0.014$
$\checkmark$ Static $0.680\pm 0.010$ $0.678\pm 0.012$ $0.470\pm 0.009$
$\checkmark$ Pure $0.670\pm 0.010$ $0.670\pm 0.019$ $0.468\pm 0.013$
Figure 6: The x-axis of the figure indicates the decoding technique used in the model listed in Table 4. All the decoding techniques improve the localization performance $L$. However, the classification performance $L_{\mathrm{top10}}$ does not improve accordingly, indicating an increase in false positives.
Cases Visualization
Finally, to better understand which test examples are easy and hard for our models, we visualized a few exams along with their classification scores. Figures 8 and 9 showcase images where the model assigned cancerous objects high malignant scores (scores $\geq 0.8$) and low scores (scores $\leq 0.1$), respectively. We observed that the model correctly localizes abnormal objects in all images. However, it tends to assign high scores to high-density masses featuring non-circumscribed, irregular, or indistinct borders, which are typically indicative of malignancy to the human eye. Conversely, the model usually assigns low scores to low-density masses with circumscribed borders, which can easily be confused with benign cases (Lee et al., 2018).
Figure 7: Training and validation loss of DETR models. The left panel depicts the classification loss and the right panel depicts the localization loss. Each line represents the average of five models with different random seeds. It clear that the models without any decoding techniques (the blue line) overfit significantly less to the training data than other models with decoding techniques (other color lines).
(a) $0.814$
(b) $0.823$
(c) $0.825$
(d) $0.822$
Figure 8: Example mammograms with high classification scores They tend to be higher density masses with non-circumscribed, irregular or indistinct borders, strongly suggestive of malignancy. The red bounding boxes are ground truth annotations and the yellow bounding boxes are the prediction of our model. The prediction results are provided by Deformable DETR with pure query initialization method and IBBF, as detailed in Table 4.
(a) $0.014$
(b) $0.018$
(c) $0.040$
(d) $0.044$
Figure 9: Example mammograms with lowest classification scores They are mostly asymmetric tissue or low density masses with circumscribed borders, more likely to be false-positives. The red bounding boxes are ground truth annotations and the yellow bounding boxes are the prediction of our model. The prediction results are provided by Deformable DETR with pure query initialization method and IBBF, as detailed in Table 4.
5 Conclusions
In this study, we investigated the impact of common design choices in Deformable DETR on medical images object detection, focusing on the NYU Breast Cancer Screening Dataset. We found out that all the design choices we experimented with need to be reconsidered and a simpler architectures typically lead to better performance on medical dataset. Our results further conclude that while the model effectively identifies abnormal tissues, it struggles to distinguish benign from malignant lesions and assigning high classification scores to the latter. This issue partly stems from the relatively small proportion of positive cases, as only about $10\%$ of the images contain positive objects. Partly because visual features typically used for classification in natural image datasets, such as object size, shape, and color, are less discriminative in medical images. For example malignant lesions in our images often lack distinct borders, have uniform sizes, and are presented in grayscale.
Future research should prioritize developing techniques for extracting subtle features relevant to medical imaging, including texture variations, density changes, irregular borders, and microcalcifications. Additionally, exploring architectural designs capable of efficiently processing full-resolution images is crucial, enabling models to focus on rich information in the pertinent regions while disregarding extensive background areas. Additionally, implementing regularization methods to mitigate overfitting in classification tasks, especially in datasets with limited positive cases, is vital for improving model performance.
6 Limitations and future work
Our study had several limitations. First, while we focused on Deformable DETR, comparing its performance with other transformer-based object detection architectures, such as DINO, would provide a broader understanding of the strengths and weaknesses of different models. Future studies should explore multiple architectures to reach more generalizable conclusions. Additionally, due to the amount of the resources necessary to conduct the experiments with a large medical dataset, our investigation was confined to a single dataset. To ensure the robustness of our findings, it is important to validate these results across various medical imaging datasets. Considering the vast diversity within medical image datasets, including modalities such as brain MRI or ultrasound images, future research should encompass these distinct characteristics to provide a comprehensive evaluation.
Acknowledgments
This work was supported in part by a grants from the National Institutes of Health (P41EB017183), the National Science Foundation (1922658), the Gordon and Betty Moore Foundation (9683), and the Mary Kay Ash Foundation (05-22). We also gratefully acknowledge the support of Nvidia Corporation with the donation of some of the GPUs used in this research.
Ethical Standards
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
Conflicts of Interest
The authors do not declare any conflicts of interest.
Data availability
Our internal (NYU Langone Health) dataset is not publicly available due to internal data transfer policies. We released a data report on data curation and preprocessing to encourage reproducibility. The data report can be accessed at this link.
A.1 Multi-head self-attention (MHSA)
A standard MHSA with $M$ heads is defined as:
$MHSA(Q,K,V)=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(\frac{QW_{mq}(KW_{mk})^{T}}{\sqrt{d/M}}\right)VW_{mv}\right].$ (3)
The $K$, $Q$ and $V$ represent the query, key, and value matrices respectively, defined with respect to input feature map $x_{f}\in\mathbb{R}^{WH\times d}$ and its positional embedding $x_{p}\in\mathbb{R}^{WH\times d}$:
$Q=x_{f}+x_{p},K=x_{f}+x_{p},V=x_{f}.$ (4)
$W_{mq},W_{mk},W_{mv}\in\mathbb{R}^{d\times d/M}$ linearly transforms $K,Q,V$ in $m$-th head and $W_{mo}\in\mathbb{R}^{d/M\times d}$.
A.2 Multi-head (MH) cross-attention
The MH cross-attention module performs the same computation as MHSA defined in 3, except that $K,Q,V$ are defined based on two different set of tokens. The queries $Q$ is defined by object queries $q=q_{c}+q_{p}$ where $q_{p}$ and $q_{c}$ are the positional embedding and content embedding of the object queries. The keys $K$ are defined by encoder features $x_{enc}+x_{p}$. Specifically,
$Q=q_{c}+q_{p},K=x_{enc}+x_{p},V=q_{c}.$ (5)
A.3 Set prediction loss
DETR uses a set prediction loss which enables end-to-end training without non-maximum suppression (NMS). DETR produces a fixed number of predictions per image $N$, and $N$ is set to be significantly larger than the maximum possible number of objects in the image. Let $\{\hat{y}_{i}=(\hat{c}_{i},\hat{b}_{i})\}_{N}$ be all pairs of class and box predictions. The set of $N$ labels is $\{y_{i}=(c_{i},b_{i})\}_{N}$ where each ground truth label represents an object in the image. If there are fewer objects than $N$, the rest of the labels are the empty classes $(0,\emptyset)$. The set prediction loss is computed in two steps. The first step is to find a permutation $\sigma$ on the set of labels $\{y_{i}\}$ that minimizes the matching loss, defined as below:
$\mathbf{\hat{\sigma}}=\arg\min_{\sigma\in\Sigma_{N}}\sum_{N}^{i}\mathcal{L}_{match}(\hat{y}_{i},y_{\sigma(i)}).$
The matching loss for a matching pair is a linear combination of classification loss, box regression loss and GIoU loss (Rezatofighi et al., 2019). The classification loss is a standard focal loss (Lin et al., 2017). The regression loss and the GIoU loss are only applied to non-empty labels. It is defined as the following:
$\mathcal{L}_{match}(\hat{y}_{i},y_{\sigma(i)})=W_{cls}\mathcal{L}_{cls}(\hat{c}_{i},c_{\sigma(i)})+1_{b\neq\emptyset}\left(W_{l1}\mathcal{L}_{l1}(\hat{b}_{i},b_{\sigma(i)})+W_{giou}\mathcal{L}_{giou}(\hat{b}_{i},b_{\sigma(i)})\right),$
where $W_{cls},W_{l1},W_{giou}$ are scalar coefficients that are tuned as hyperparameters to balance the scale of different losses. The Hungarian algorithm (Kuhn, 1955) can efficiently find the optimal matching $\hat{\sigma}$. The second step is to minimize the loss function $\sum_{N}^{i}\mathcal{L}_{match}(\hat{y}_{i},y_{\hat{\sigma}(i)})$ with the permutation $\hat{\sigma}$ on the label set.
DETR also utilizes auxiliary loss in each decoder layer to provide stronger supervision. At the end of each decoder layer, it predicts $N$ boxes and class scores with MLP prediction heads. All the prediction heads share weights. The above two steps, the matching step and Hungarian loss minimization, are applied to each decoder layer’s output. At inference, only the last layer’s output are used as the final prediction.
B Deformable DETR architecture
B.1 Deformable multi-head self-attention
Formally, deformable MHSA for a single query $q\in\mathbb{R}^{d}$ in the feature map is given by:
$\mathrm{Deformable\_MHSA}(Q_{q},K,V)=\sum_{m=1}^{M}W_{mo}\left[\mathrm{softmax}\left(K_{q}W_{mk}\right)V_{q}W_{mv}\right]$ (6)
The $Q$, $K$ and $V$ represent the query, key, and value matrices respectively, defined as the following,
$Q=x_{f}+x_{p}\text{ and }Q_{q}=q,K_{q}=\delta(K,q)\in\mathbb{R}^{k\times d},V_{q}=\delta(V,q)\in\mathbb{R}^{k\times d}.$ (7)
The key sampling function $\delta$ samples $k$ keys from the full set of keys $K=x_{f}+x_{p}$ by generating the sampling offsets $\Delta p$ with respect to reference points $p_{q}$: $\delta(K,q)=K(p_{q}+\Delta p)$. The sampling offsets are obtained through a linear transformation of the query $q$.
C Iterative Bounding Box Refinement Technique
In the standard Deformable DETR, a 2D reference point $r_{q}\in[0,1]^{2}$ for each object query $q$ is derived from its learnable positional embedding $p_{q}$ via a linear layer
$r_{q}=\mathrm{linear}(p_{q}).$
Throughout the decoder, the locations of these reference points remain constant. They are updated based on the learnable positional embedding $p_{q}$ when a backward pass is completed. Formally, let $r_{q}^{i}$ be the reference points of an object query $q$ in i-th decoder layer. In standard Deformable DETR,
$r_{q}^{1}=r_{q}^{2}=\ldots=r_{q}^{6}.$
In IBBR, the reference points $r_{q}^{i}$ in i-th decoder layer is refined based on the previous reference points $r_{q}^{i-1}$ and the offsets predicted by auxiliary prediction head.
$r_{q}^{i}=S(\mathrm{linear}^{-1}(S^{-1}(r_{q}^{i-1}))+\mathrm{MLP}(x_{dec}^{i})),$
where $S$ and $S^{-1}$ represent the sigmoid function and its inverse, and $x^{i}_{dec}$ is the output of the i-th decoder layer.
D Backbone Pre-training
We pretrained the Swin-T Transfromer backbone with a cancer classification task on our dataset. This classification task is a binary multi-label classification that predicts two scores indicating if an input image contains benign lesions or malignant lesions.
E Hyperparameter Tuning
Our method for hyper-parameter tuning is random search. We tuned the following hyperparameters and their ranges on quarter resolution images:
  • learning rate $\eta\in 10^{[3,5.5]}$,
  • scale of the backbone learning rate $s\in[1,0.01]$ (backbone learning rate = $s\times\eta$ ),
  • weight decay $\lambda\in 10^{[3,6]}$,
  • number of object queries $N\in[10,200]$,
  • two hyperparameters $\alpha$ and $\gamma$ in the focal loss $\alpha\in[0,1]$, $\gamma\in[0,3]$,
  • the coefficients on classification loss and GIoU loss $\in[0,1]$.
We train $80$ jobs in total and choose the best model based on FAUC_1.